The table describes hypothetical employment statistics for a country in a given year. Full-time workers and Labor force Involuntary part-time workers Discouraged workers voluntary part-time workers 188 milliorn 200 milliorn 6 million 4 million Based on the information, what is the unemployment rate? Enter your answer as percentage rounded to one decimal place.
Answer #1

Labor force = Employed + Unemployed

Employed = (188+6)=194 million

Labor force = 200 million

Unemployed = (200-194) million = 6 million

Unemployment rate = (6/200)100 = 3%